A printing company's bookbinding machine has a probability of 0.75 for producing a perfect book. This machine used to bind three books together.
1) Find the probability that all three books are perfect books?

2) Find the probability that at least one of the books are perfect?

Answer by stanbon(75887)

A printing company's bookbinding machine has a probability of 0.75 for producing a perfect book. This machine used to bind three books together.
1) Find the probability that all three books are perfect books?
----
Binomial Prob. with n = 3 and P(perfect( = 0.75
P(x = 3) = 0.75^3 = 0.4219
--------------------------------
2) Find the probability that at least one of the books are perfect?
P(at least one perfect) = 1 - P(all are defective)
----
= 1 - 0.25^3
----
= 0.9844
============
Cheers,
Stan H.
============